OpenStudy (anonymous):

If 0^2/3 = 0, why is 0^-2/3 undefined?

OpenStudy (anonymous):

Because anything raised to the negative power becomes its reciprocal, here lemme show you what i mean: 2^-1 = 1/2 3^-1 = 1/3 x^-1 = 1/x so 0 taken to a negative power is just 1/0 = ??? I don't know about you, but I never learned to divide by zero, at least not without calculus ;) so let's rewrite your problem as: 0^-2/3 = 1/0^(2/3) = 1/0 = undefined

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (anonymous):

0/3 = 0 is totally different with: 0^-2/3 = 1/0 * 3 = 1/0 ( undefined!)